# Visual Cue Enhancement and Dual Low-Rank Adaptation for Efficient Visual Instruction Fine-Tuning

### v1

@@ Abstract
Fine-tuning multimodal large language models (MLLMs) presents significant challenges, including a reliance on high-level visual features that limits fine-grained detail comprehension, and data conflicts that arise from task complexity. To address these issues, we propose an efficient fine-tuning framework with two novel approaches: Vision Cue Enhancement (VCE) and Dual Low-Rank Adaptation (Dual-LoRA). VCE enhances the vision projector by integrating multi-level visual cues, improving the model’s ability to capture fine-grained visual features. Dual-LoRA introduces a dual low-rank structure for instruction tuning, decoupling learning into skill and task spaces to enable precise control and efficient adaptation across diverse tasks. Our method simplifies implementation, enhances visual comprehension, and improves adaptability. Experiments on both downstream tasks and general benchmarks demonstrate the effectiveness of our proposed approach.
@@ 1 Introduction
Despite the great success, existing MLLMs have two main limitations. On the one hand, in the vision projector pretraining stage, current methods primarily rely on high-level visual features, often overlooking low-level and fine-grained details, which limits visual comprehension. For example, most works[15, 2] use only the high-level semantic feature map for vision token projection, while some [24, 31] leverage multi-layer visual features yet still miss finer details. On the other hand, in the instruction-tuning stage, as downstream tasks grow in diversity and complexity, data conflicts in LoRA-based instruction tuning become more prominent  [4, 9]. To mitigate this, recent studies [4, 18] incorporate the Mixture of Experts (MoE) paradigm into LoRA modules, embedding multiple LoRA units within a single linear layer so that each unit can leverage its specific strengths. To better accommodate varied tasks, LoRA experts with different levels of granularity or customized routing strategies [18] have been introduced. However, this complex design not only increases implementation complexity but may also extend the time required for training and inference. We thus raise the question: Is there a simple and unified approach to flexibly represent any group of LoRA experts?
@@
Experiments on downstream tasks and general MLLM benchmarks demonstrate the effectiveness of our proposed method. Our contributions can be summarized as follows:
@@
Figure 2: Illustration of our proposed components. The Visual Cue Enhancement (VCE) module uses the projected vision feature map $f^{\textbf{a}}$ selected during pretraining as the anchor feature, leveraging nearby patches in multi-layer feature maps $[f^{\textbf{ref}}]$ to enrich the local visual cues of $f^{\textbf{a}}$. The Dual Low-Rank Adaptation decomposes the input feature into skill space and task space via $S$ and $T$, which are then multiplied element-wise, projected back to the same dimension as $x$ via $B$, added to the features obtained from the pretrained weights $W$, and passed to the next layer.
@@ 3.3 Multi-level Local Visual Cue Enhancement
The enhanced vision token $t_{\text{vision}}$ is then concatenated with text tokens and fed into the multimodal model $G$ for further processing.
@@ 3.4 Dual Low Rank Adaptation
Corollary 2: For a LoRA $BA$ with rank $r$, the LoRA $B(A\cdot\sigma(T))$ can be decomposed into any set of LoRA groups $\{B_{k}A_{k}\}_{k=1}^{K}$, provided that the constraint $r\geq\sum_{k=1}^{K}\text{rank}(B_{k}A_{k})$ holds. Here, $T$ is a matrix with the same shape as $A$, and $\sigma$ is a non-linear activation function.
@@ 3.4.1 Skill Low-Rank Space
In practice, however, a split LoRA expert group often outperforms a single, larger LoRA [10]. This can be attributed to variations in routing strategies and initialization across individual LoRAs. For example, LLaVA-MoLE [4] activates the top-k experts for targeted learning, while RoDE [10] employs heterogeneous LoRA experts, each initialized with distinct distributions. The complex training process, particularly the routing score, leads to divergence in the representational spaces learned by different LoRAs over time. The complex training process, especially the routing score, leads to divergence in the representational spaces learned by different LoRAs over time. Intuitively, even for the same input, each expert may experience unique gradients due to its intrinsic numerical differences and routing scores.
@@ 3.4.2 Rectified Skill Low-Rank Space
$D(x)$ $=\frac{r}{\alpha}B(Sx\cdot\sigma(Tx)),$ (9)
@@
$D(x)$ $=\frac{r}{\alpha}B(\text{Norm}(Sx)\cdot\sigma(Tx)).$ (10)
@@
$z$ $=W(x)+\frac{r}{\alpha}B(\text{Norm}(Sx)\cdot\sigma(Tx)).$ (11)
@@ 4.1 Model and Optimizer Configuration
Our VCE module requires only 5.52 MB of memory, compared to the 40.02 MB needed for the vision projector in LLaVA-1.5-7B. Furthermore, our Dual-LoRA task consumes approximately 1.5 MB per LoRA rank when injected solely into the query projection and value projection layers. For optimization, we use the Adam optimizer across all trainable parameters. During the pretraining stage, we set a global batch size of 256 and a learning rate of 0.001 with no weight decay. In the fine-tuning stage, we use a global batch size of 128, setting the learning rate to 0.0002 for the LMM adapter and 0.00002 for the vision feature projector, respectively.
@@ 4.2.1 Experiment Setup
Baselines. We select several Mixture of Experts (MoE) methods within the LoRA framework: a sparse activation method, MoE-LoRA (top-2), which activates the top-2 experts; a dense activation method, MoE-LoRA (softmax), which uses a softmax router to engage all experts; and a rectified activation method, RoDE [10], with rectified diverse expert activation. Both MoE-LoRA (top-2) and MoE-LoRA (softmax) are configured with ranks $[16,16,16,16]$, while RoDE is set with varying ranks $[32,16,8,8]$. Additionally, we include a standard LoRA baseline with rank 64, and our Dual-LoRA is also set to rank 64.
Implementation Details. We conduct our experiments on a setup with 8 RTX 4090 GPUs. The LLM adapters are integrated only within the linear layers of the query and value projectors in the transformer blocks, allowing us to train the model on a single RTX 4090 GPU if necessary.

### v2

@@ Abstract
Parameter-efficient fine-tuning multimodal large language models (MLLMs) presents significant challenges, including a reliance on high-level visual features that limits fine-grained detail comprehension, and data conflicts that arise from task complexity. To address these issues, we propose an efficient fine-tuning framework with two novel approaches: Vision Cue Enhancement (VCE) and Dual Low-Rank Adaptation (Dual-LoRA). VCE enhances the vision projector by integrating multi-level visual cues, improving the model’s ability to capture fine-grained visual features. Dual-LoRA introduces a dual low-rank structure for instruction tuning, decoupling learning into skill and task spaces to enable precise control and efficient adaptation across diverse tasks. Our method simplifies implementation, enhances visual comprehension, and improves adaptability. Experiments on both downstream tasks and general benchmarks demonstrate the effectiveness of our proposed approach.
@@ 1 Introduction
Despite the great success, existing MLLMs have two main limitations. On the one hand, in the vision projector pretraining stage, current methods primarily rely on high-level visual features, often overlooking low-level and fine-grained details, which limits visual comprehension. For example, most works[15, 2] use only the high-level semantic feature map for vision token projection, while some [24, 31] leverage multi-layer visual features yet still miss finer details. On the other hand, in the instruction-tuning stage, as downstream tasks grow in diversity and complexity, data conflicts in LoRA-based instruction tuning become more prominent  [4, 9]. To mitigate this, recent studies [4, 18] incorporate the Mixture of Experts (MoE) paradigm into LoRA modules, embedding multiple LoRA units within a single linear layer so that each unit can leverage its specific strengths. To better accommodate varied tasks, LoRA experts with different levels of granularity [10] or customized routing strategies [18] have been introduced. However, this complex design not only increases implementation complexity but may also extend the time required for training and inference. We thus raise the question: Is there a simple and unified approach to flexibly represent any group of LoRA experts?
@@
Experiments conducted on downstream tasks within the food domain, as well as on several general MLLM benchmarks, highlight the effectiveness of our proposed method. Our contributions can be summarized as follows:
@@
Figure 2: Illustration of our proposed components. The Visual Cue Enhancement (VCE) module uses the projected vision feature map $f^{\textbf{a}}$ selected during pretraining as the anchor feature, leveraging nearby patches in multi-layer feature maps $[f^{\textbf{ref}}]$ to enrich the local visual cues of $f^{\textbf{a}}$. The Dual Low-Rank Adaptation (Dual-LoRA) decomposes the input feature into skill space and task space via $S$ and $T$, which are then multiplied element-wise, projected back to the same dimension as $x$ via $B$, added to the features obtained from the pretrained weights $W$, and passed to the next layer.
@@ 3.3 Multi-level Local Visual Cue Enhancement
The enhanced vision token $t_{\text{vision}}$ is then concatenated with text tokens and fed into the large language modal model $G$ for further processing. The structure of the adaptive $G$ is detailed in the next section.
@@ 3.4 Dual Low Rank Adaptation
Corollary 2: For a LoRA $BA$ with rank $r$, the LoRA $B(A\odot\sigma(T))$ can be decomposed into any set of LoRA groups $\{B_{k}A_{k}\}_{k=1}^{K}$, provided that the constraint $r\geq\sum_{k=1}^{K}\text{rank}(B_{k}A_{k})$ holds. Here, $T$ is a matrix with the same shape as $A$, and $\sigma$ is a non-linear activation function.
@@ 3.4.1 Skill Low-Rank Space
In practice, however, a split LoRA expert group often outperforms a single, larger LoRA [10]. This can be attributed to variations in routing strategies and initialization across individual LoRAs. For example, LLaVA-MoLE [4] activates the top-k experts for targeted learning, while RoDE [10] employs heterogeneous LoRA experts, each initialized with distinct distributions. The complex training process, particularly the routing score, leads to divergence in the representational spaces learned by different LoRAs over time. Intuitively, even for the same input, each expert may experience unique gradients due to its intrinsic numerical differences and routing scores.
@@ 3.4.2 Rectified Skill Low-Rank Space
$D(x)$ $=\frac{r}{\alpha}B(Sx\odot\sigma(Tx)),$ (9)
@@
$D(x)$ $=\frac{r}{\alpha}B(\text{Norm}(Sx)\odot\sigma(Tx)).$ (10)
@@
$z$ $=Wx+\frac{r}{\alpha}B(\text{Norm}(Sx)\odot\sigma(Tx)).$ (11)
@@ 4.1 Model and Optimizer Configuration
Our VCE module requires only 5.52 MB of memory, compared to the 40.02 MB needed for the vision projector in LLaVA-1.5-7B. Furthermore, our Dual-LoRA task consumes approximately 1.5 MB per LoRA rank when injected solely into the query projection and value projection layers. For optimization, we use the Adam optimizer across all trainable parameters. During the pretraining stage, we set a global batch size of 256 and a learning rate of 0.001 with no weight decay. In the fine-tuning stage, we use a global batch size of 128, setting the learning rate to 0.0002 for the LLM adapter and 0.00002 for the vision feature projector, respectively.
@@ 4.2.1 Experiment Setup
Baselines. We select several Mixture of Experts (MoE) methods within the LoRA framework: a sparse activation method, MoE-LoRA (top-2), which activates the top-2 experts; a dense activation method, MoE-LoRA (softmax), which uses a softmax router to engage all experts; and a rectified activation method, RoDE [10], with rectified diverse expert activation. Both MoE-LoRA (top-2) and MoE-LoRA (softmax) are configured with ranks $[16,16,16,16]$, while RoDE is set with varying ranks $[32,16,8,8]$. Additionally, we include a vanilla LoRA baseline with rank 64, and our Dual-LoRA is also set to rank 64.
Implementation Details. We conduct our experiments on a setup with 8 RTX 4090 GPUs. The LLM adapters are integrated only within the linear layers of the query and value projectors in the LLM transformer blocks, allowing us to train the model on a single RTX 4090 GPU if necessary.